OpenStudy (anonymous):

show tat√((1-x)/(1+x))=1-x+(1/2)(x^2)

OpenStudy (thomas5267):

Is it \(\sqrt{\frac{1-x}{1+x}}=1-x+\frac{1}{2x^2}\)?

OpenStudy (anonymous):

no...is (1/2)(x^2)

OpenStudy (thomas5267):

\(\sqrt{\frac{1-x}{1+x}}=1-x+\frac{1}{2}x^2\)?

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 })x ^{2}\]

OpenStudy (shamil98):

just substitute x with any number to see if it is true..